# Consistency Models Made Easy

### v1

@@ Appendix B Exploring the Design Space of Consistency Models
Appendix C Experimental Details
@@
Model Setups Uncond CIFAR-10 Cls-Cond CIFAR-10 ImageNet 64$\times$64 ECM-S ECM-M ECM-L ECM-XL Model Channels $128$ $128$ $192$ $256$ $320$ $384$ Model capacity (Mparams) $55.7$ $55.7$ $280.2$ $497.8$ $777.5$ $1119.3$ Model complexity (GFLOPs) $21.3$ $21.3$ $101.9$ $180.8$ $282.2$ $405.9$ Training Details Training Duration (Mimg) $12.8$ $12.8$ $12.8$ $12.8$ $12.8$ $12.8$ Minibatch size $128$ $128$ $128$ $128$ $128$ $128$ Iterations $100$k $100$k $100$k $100$k $100$k $100$k Dropout probability $20\%$ $20\%$ $40\%$ $50\%$ $50\%$ $50\%$ Optimizer RAdam RAdam Adam Adam Adam Adam Learning rate max ($\alpha_{\text{ref}}$) $0.0001$ $0.0001$ $0.0010$ $0.0090$ $0.0080$ $0.0070$ Learning rate decay ($t_{\text{ref}}$) - - $2000$ $2000$ $2000$ $2000$ EMA beta $0.9999$ $0.9999$ Training Cost Number of GPUs $1$ $1$ $4$ $8$ $8$ $8$ GPU types A6000 A6000 H100 H100 H100 H100 Training time (hours) $24$ $24$ $8.5$ $8.5$ $12$ $15$ Generative Performance $1$-step FID $4.54$ $3.81$ $5.51$ $3.67$ $3.55$ $3.35$ $2$-step FID $2.20$ $2.02$ $3.18$ $2.35$ $2.14$ $1.96$ ECT Details Regular Weighting ($\bar{w}(t)$) $\nicefrac{{1}}{{(t-r)}}$ $\nicefrac{{1}}{{(t-r)}}$ $\nicefrac{{1}}{{t^{2}}}+\nicefrac{{1}}{{\sigma_{\text{data}}^{2}}}$ $\nicefrac{{1}}{{t^{2}}}+\nicefrac{{1}}{{\sigma_{\text{data}}^{2}}}$ $\nicefrac{{1}}{{t^{2}}}+\nicefrac{{1}}{{\sigma_{\text{data}}^{2}}}$ $\nicefrac{{1}}{{t^{2}}}+\nicefrac{{1}}{{\sigma_{\text{data}}^{2}}}$ Adaptive Weighting ($w(\Delta)$) Adaptive Weighting Smoothing ($c$) $0.0$ $0.0$ $0.06$ $0.06$ $0.06$ $0.06$ Noise distribution mean ($P_{\text{mean}}$) $-1.1$ $-1.1$ $-0.8$ $-0.8$ $-0.8$ $-0.8$ Noise distribution std ($P_{\text{std}}$) $2.0$ $2.0$ $1.6$ $1.6$ $1.6$ $1.6$
Model Setup.
For both unconditional and class-conditional CIFAR-10 experiments, we initial ECMs from the pretrained EDM [26] of DDPM++ architecture [66]. For class-conditional ImageNet 64$\times$64 experiments, we initial ECM-S/M/L/XL, ranging from $280\mathrm{M}$ to $1.1\mathrm{B}$, from the pretrained EDM2 [28]. Detailed model configurations are presented in Tab. 6.
We follow Karras et al. [26], Song et al. [67] and set $c_{\text{skip}}(t)=\nicefrac{{\sigma_{\text{data}}^{2}}}{{(t^{2}+\sigma_{\text{data}}^{2})}}$ and $c_{\text{out}}(t)=\nicefrac{{t\sigma_{\text{data}}}}{{\sqrt{t^{2}+\sigma_{\text{data}}^{2}}}}$, where $\sigma_{\text{data}}^{2}$ is the variance of (normalized) data, and set to $0.5$ for both CIFAR-10 and ImageNet 64$\times$64.
Computational Cost.
ECT is computationally efficient. On ImageNet 64$\times$64, ECT requiring only $\mathbf{0.39\%}$ of the iCT [64] training budget, and $\mathbf{0.60\%}$ to $\mathbf{1.91\%}$ of the EDM2 [28] pretraining budget depending on the model sizes. The exact computational resources required to train each individual model is shown in Tab. 6.
@@
We use RAdam [39] optimizer for experiments on CIFAR-10 and Adam [31] optimizer for experiments on ImageNet 64$\times$64. We set the $\beta$ to ($0.9$, $0.999$) for CIFAR-10 and ($0.9$, $0.99$) for ImageNet 64$\times$64. All the hyperparameters are indicated in Tab. 6. We do not use any learning rate decay, weight decay, or warmup on CIFAR-10. We follow EDM2 [28] to apply an inverse square root learning rate decay schedule on ImageNet 64$\times$64.
On CIFAR-10, we employ the traditional Exponential Moving Average (EMA). To better understand the influence of the EMA rate, we track three Power function EMA models on ImageNet 64$\times$64, using EMA lengths of $0.01$, $0.05$, and $0.10$. The multiple EMA models introduce no visible cost to the training speed. Considering our training budget is much smaller than the diffusion pretraining stage, we didn’t perform the Post-Hoc EMA search as in EDM2 [28].
Experiments for ECT are organized in a non-adversarial setup to better focus and understand CMs and avoid inflated FID [68]. We conducted ECT using full parameter tuning in this work, even for models over $1\mathrm{B}$ parameters. Investigating the potential of Parameter Efficient Fine Tuning (PEFT) [24] can further reduce the cost of ECT to democratize efficient generative models, which is left for future research.
We train multiple ECMs with different choices of batch sizes and training iterations. By default, ECT utilizes a batch size of $128$ and $100$k iterations, leading to a training budget of $12.8\mathrm{M}$. We have individually indicated other training budgets alongside the relevant experiments, wherever applicable.
@@
We apply stochastic sampling for 2-step generation. For 2-step sampling, we follow Song and Dhariwal [64] and set the intermediate $t=0.821$ for CIFAR-10, and $t=1.526$ for ImageNet 64$\times$64.
Evaluation Details.
For both CIFAR-10 and ImageNet 64$\times$64, FID and $\text{FD}_{\text{DINOv2}}$ are computed using $50$k images sampled from ECMs. As suggested by recent works [68, 28], $\text{FD}_{\text{DINOv2}}$ aligns better with human evaluation. We use dgm-eval [^1] https://github.com/layer6ai-labs/dgm-evalto calculate $\text{FD}_{\text{DINOv2}}$ [68] to ensure align with previous practice.
@@
Scaling Laws of Training Compute.
@@
The first two points of $2^{0}$ and $2^{1}$ on Fig. 4 (Left) use a batch size of $128$ for $100$k and $200$k iterations, respectively. The third point of $2^{2}$ corresponds to ECM trained with batch sizes of $256$ for $200$k iterations. The final three points of $2^{3}$, $2^{4}$, and $2^{5}$ correspond to ECM trained with a batch size of $512$ for $200$k, $400$k, and $800$k iterations, respectively, with the smoothing factor $c=0.03$ enabled in the adaptive weighting $w(\Delta)$. We use $\bar{w}(t)=\nicefrac{{1}}{{t}}$ as the regular weighting function to train all these models as this $\bar{w}(t)$ achieves good performance on $\text{FD}_{\text{DINOv2}}$.
Scaling Laws of Model Size and Model FLOPs.
We have included details of model capacity as well as FLOPs in Tab. 6 to replicate this plot on ImageNet 64$\times$64.
On ImageNet 64$\times$64, we scale up the training budgets of ECM-S and ECM-XL from $12.8\mathrm{M}$ (batch size of $128$ and $100$k iterations) to $102.4\mathrm{M}$ (batch size of $1024$ and $100$k iterations). We empirically find that scaling the base learning rate by $\sqrt{n}$ works well when scaling the batch size by a factor of $n$ on ImageNet 64$\times$64 ECMs when using Adam [31] optimizer.
Appendix D Broader Impacts

### v2

@@ Appendix B Exploring Design Space & Scaling of Consistency Models
Appendix C Extension to Consistency Distillation.
Continuous-time Consistency Distillation
Consistency Distillation (CD) (Song et al., 2023) uses a fixed discrete schedule derived from a specific sampler, such as the EDM sampler. This approach inherently limits $\Delta t\to{\mathrm{d}}t$, therefore causing a non-trivial discretization error on approximating the differential consistency condition. Building upon ECT, we extend our continuous-time schedule to Consistency Distillation, which we term Easy Consistency Distillation (ECD).
On ImageNet 64$\times$64, we implement Consistency Distillation as the baseline using pretrained EDM2-S (Karras et al., 2024) with weighting functions, noise distribution, and dropout. The results, shown in Tab. 7, demonstrate that continuous-time distillation improves upon the standard CD approach.
Given well-pretrained DMs, ECD typically demonstrates a performance advantage over ECT at smaller batch sizes (e.g., 128), primarily due to the variance reduction in approximating the score function:
$\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t})=\mathbb{E}[\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t}|{\mathbf{x}}_{0})|{\mathbf{x}}_{t}]$
This advantage stems from the teacher models, which provide a variance-reduced estimate of the score function compared to the Monte Carlo estimation used in ECT. However, this gap tends to narrow as we scale up to larger model sizes. Scaling compensates for the limited training budget and variances, closing the gap between the two.
Table 7: Easy Consistency Distillation (ECD) on ImageNet 64$\times$64 using the same budget of $12.8\mathrm{M}$ training images (batch size 128 and 100k iterations) as ECT in Tab. 1.
Methods 1-step FID 2-step FID
CD-S 8.18 3.71
ECD-S 3.33 2.10
ECD-M 2.78 1.92
ECD-XL 2.54 1.77
Consequently, while ECD may outperform ECT in resource-constrained scenarios or with smaller models, the distinction becomes less pronounced as we move to larger scales. This observation underscores the scaling factor (computational resources and model size) when choosing between ECT and ECD for a given application.
Data-Free ECD.
Inspired by the recent progress on data-free distillation (Gu et al., 2023, Luo et al., 2024, Yin et al., 2023, Zhou et al., 2024a), we explore a data-free variant of Easy Consistency Distillation (ECD). Instead of sampling data points ${\mathbf{x}}_{0}$ from a given dataset $\mathcal{D}$, we generate synthetic data points directly from the CM itself on the fly, ${\mathbf{x}}_{0}=f_{\operatorname{sg}(\theta)}(\bm{\epsilon}^{\prime},T)$, where $\bm{\epsilon}\textquoteright\sim p(\bm{\epsilon})$, $T$ is the maximum noise level for sampling, followed by the same ECD training step over the self synthetic data. This eliminates the need for a distillation dataset or the construction of synthetic data from the teacher model (Luhman and Luhman, 2021, Geng et al., 2024, Zheng et al., 2023).
On ImageNet 64$\times$64, this data-free ECD achieves 1-step FID of 4.38 and 2-step FID of 2.77, comparable to the data-dependent ECD and ECT schemes at the same budget. It suggests that data-free ECD can be a competitive alternative in scenarios where access to large datasets is limited or unavailable. We summarize both ECD and its data-free variant in Alg. 2.
Algorithm 2 Easy Consistency Distillation (ECD) / Data-Free ECD
Input: Dataset ${\mathcal{D}}$ (for ECD), a pretrained diffusion model $\phi$, mapping function $p(r\mid t,\mathrm{Iters})$, weighting function $w(t)$.
Init: $\theta\leftarrow\phi$, $\mathrm{Iters}=0$.
repeat
     if ECD then
         $\mathrm{Sample}$ ${\mathbf{x}}_{0}\sim\mathcal{D}$
     else if Data-Free ECD then
         $\mathrm{Sample}$ $\bm{\epsilon}^{\prime}\sim p(\bm{\epsilon})$
         $\mathrm{Compute}$ ${\mathbf{x}}_{0}=f_{\operatorname{sg}(\theta)}(\bm{\epsilon}^{\prime},T)$triangleright $\mathrm{sg}$ is stop-gradient operator
     end if
     $\mathrm{Sample}$ $\bm{\epsilon}\sim p(\bm{\epsilon})$, $t\sim p(t)$, $r\sim p(r\mid t,\mathrm{Iters})$
     $\mathrm{Compute}$ ${\mathbf{x}}_{t}={\mathbf{x}}_{0}+t\cdot\bm{\epsilon}$, $\Delta t=t-r$
     $\mathrm{Compute}$ ${\mathbf{x}}_{r}=\text{Solver}(\mathbf{x}_{t},t\to r,f_{\phi})$
     $L(\theta)=w(t)\cdot{\mathbf{d}}(f_{\theta}({\mathbf{x}}_{t}),f_{\operatorname{sg}(\theta)}({\mathbf{x}}_{r}))$
     $\theta\leftarrow\theta-\eta\nabla_{\theta}L(\theta)$
     $\mathrm{Iters}=\mathrm{Iters}+1$
until $\Delta t\to\mathrm{d}t$ return $\theta$triangleright ECD
Appendix D Experimental Details
Model Setup.
For both unconditional and class-conditional CIFAR-10 experiments, we initial ECMs from the pretrained EDM (Karras et al., 2022) of DDPM++ architecture (Song et al., 2021b). For class-conditional ImageNet 64$\times$64 experiments, we initial ECM-S/M/L/XL, ranging from $280\mathrm{M}$ to $1.1\mathrm{B}$, from the pretrained EDM2 (Karras et al., 2024). Detailed model configurations are presented in Tab. 8.
We follow (Karras et al., 2022, Song et al., 2023) and set $c_{\text{skip}}(t)=\nicefrac{{\sigma_{\text{data}}^{2}}}{{(t^{2}+\sigma_{\text{data}}^{2})}}$ and $c_{\text{out}}(t)=\nicefrac{{t\sigma_{\text{data}}}}{{\sqrt{t^{2}+\sigma_{\text{data}}^{2}}}}$, where $\sigma_{\text{data}}^{2}$ is the variance of (normalized) data, and set to $0.5$ for both CIFAR-10 and ImageNet 64$\times$64.
Computational Cost.
ECT is computationally efficient. On ImageNet 64$\times$64, the tuning stage of ECT requiring only $0.39\%$ of the iCT (Song and Dhariwal, 2023) training budget, and $0.60\%$ to $1.91\%$ of the EDM2 (Karras et al., 2024) pretraining budget depending on the model sizes. The exact computational resources required to train each individual model are shown in Tab. 8.
@@ Appendix D Experimental Details
Model Setups Uncond CIFAR-10 Cls-Cond CIFAR-10 ImageNet 64$\times$64 ECM-S ECM-M ECM-L ECM-XL Model Channels $128$ $128$ $192$ $256$ $320$ $384$ Model capacity (Mparams) $55.7$ $55.7$ $280.2$ $497.8$ $777.5$ $1119.3$ Model complexity (GFLOPs) $21.3$ $21.3$ $101.9$ $180.8$ $282.2$ $405.9$ Training Details Training Duration (Mimg) $12.8$ $12.8$ $12.8$ $12.8$ $12.8$ $12.8$ Minibatch size $128$ $128$ $128$ $128$ $128$ $128$ Iterations $100$k $100$k $100$k $100$k $100$k $100$k Dropout probability $20\%$ $20\%$ $40\%$ $50\%$ $50\%$ $50\%$ Dropout feature resolution - - $\leq 16\times 16$ $\leq 16\times 16$ $\leq 16\times 16$ $\leq 16\times 16$ Optimizer RAdam RAdam Adam Adam Adam Adam Learning rate max ($\alpha_{\text{ref}}$) $0.0001$ $0.0001$ $0.0010$ $0.0009$ $0.0008$ $0.0007$ Learning rate decay ($t_{\text{ref}}$) - - $2000$ $2000$ $2000$ $2000$ EMA beta $0.9999$ $0.9999$ Training Cost Number of GPUs $1$ $1$ $4$ $8$ $8$ $8$ GPU types A6000 A6000 H100 H100 H100 H100 Training time (hours) $24$ $24$ $8.5$ $8.5$ $12$ $15$ Generative Performance $1$-step FID 4.54 3.81 5.51 3.67 3.55 3.35 $2$-step FID 2.20 2.02 3.18 2.35 2.14 1.96 ECT Details Regular Weighting ($\bar{w}(t)$) $\nicefrac{{1}}{{(t-r)}}$ $\nicefrac{{1}}{{(t-r)}}$ $\nicefrac{{1}}{{t^{2}}}+\nicefrac{{1}}{{\sigma_{\text{data}}^{2}}}$ $\nicefrac{{1}}{{t^{2}}}+\nicefrac{{1}}{{\sigma_{\text{data}}^{2}}}$ $\nicefrac{{1}}{{t^{2}}}+\nicefrac{{1}}{{\sigma_{\text{data}}^{2}}}$ $\nicefrac{{1}}{{t^{2}}}+\nicefrac{{1}}{{\sigma_{\text{data}}^{2}}}$ Adaptive Weighting ($w(\Delta)$) Adaptive Weighting Smoothing ($c$) $0.0$ $0.0$ $0.06$ $0.06$ $0.06$ $0.06$ Noise distribution mean ($P_{\text{mean}}$) $-1.1$ $-1.1$ $-0.8$ $-0.8$ $-0.8$ $-0.8$ Noise distribution std ($P_{\text{std}}$) $2.0$ $2.0$ $1.6$ $1.6$ $1.6$ $1.6$
@@
We use RAdam (Liu et al., 2019) optimizer for experiments on CIFAR-10 and Adam (Kingma and Ba, 2014) optimizer for experiments on ImageNet 64$\times$64. We set the $\beta$ to ($0.9$, $0.999$) for CIFAR-10 and ($0.9$, $0.99$) for ImageNet 64$\times$64. All the hyperparameters are indicated in Tab. 8. We do not use any learning rate decay, weight decay, or warmup on CIFAR-10. We follow EDM2 (Karras et al., 2024) to apply an inverse square root learning rate decay schedule on ImageNet 64$\times$64.
On CIFAR-10, we employ the traditional Exponential Moving Average (EMA). To better understand the influence of the EMA rate, we track three Power function EMA (Karras et al., 2024) models on ImageNet 64$\times$64, using EMA lengths of $0.01$, $0.05$, and $0.10$. The multiple EMA models introduce no visible cost to the training speed. Considering our training budget is much smaller than the diffusion pretraining stage, we didn’t perform Post-Hoc EMA search as in EDM2 (Karras et al., 2024).
Experiments for ECT are organized in a non-adversarial setup to better focus and understand CMs and avoid inflated FID (Stein et al., 2024). We conducted ECT using full parameter tuning in this work, even for models over $1\mathrm{B}$ parameters. Investigating the potential of Parameter Efficient Fine Tuning (PEFT) (Hu et al., 2021) can further reduce the cost of ECT to democratize efficient generative models, which is left for future research.
We train multiple ECMs with different choices of batch sizes and training iterations. By default, experiments on ImageNet 64$\times$64 utilize a batch size of $128$ and $100$k iterations, leading to a training budget of $12.8\mathrm{M}$. We have individually indicated other training budgets alongside the relevant experiments, wherever applicable.
@@
We apply stochastic sampling for 2-step generation. For 2-step sampling, we follow (Song and Dhariwal, 2023) and set the intermediate $t=0.821$ for CIFAR-10, and $t=1.526$ for ImageNet 64$\times$64.
Intriguingly, these sampling schedules, originally developed for iCT, also perform well with our ECMs. This effectiveness across different CMs and training methods likely links to the inherent characteristics of the datasets and the forward process. Developing a scientific approach to determine optimal intermediate sampling schedules for CMs remains an open research problem.
Evaluation Metrics.
For both CIFAR-10 and ImageNet 64$\times$64, FID and $\text{FD}_{\text{DINOv2}}$ are computed using $50$k images sampled from ECMs. As suggested by recent works (Stein et al., 2024, Karras et al., 2024), $\text{FD}_{\text{DINOv2}}$ aligns better with human evaluation. We use dgm-eval [^1] https://github.com/layer6ai-labs/dgm-eval to calculate $\text{FD}_{\text{DINOv2}}$ (Stein et al., 2024) to ensure align with previous practice.
@@
1 GPU Hour Prototyping Settings.
This configuration uses a fixed $\Delta t\approx\mathrm{d}t$ by setting $q=256$ in our mapping function (corresponding to $\frac{r}{t}\approx 0.99$) and an EMA rate of 0.9993 for the model parameters. Using these settings, we run $8000$ gradient descent steps with a batch size of 128 on a single A100 40GB GPU.
Scaling of Training Compute.
@@
The first two points of $2^{0}$ and $2^{1}$ on Fig. 4 (Left) use a batch size of $128$ for $100$k and $200$k iterations, respectively. The third point of $2^{2}$ corresponds to ECM trained with batch sizes of $256$ for $200$k iterations. The final three points of $2^{3}$, $2^{4}$, and $2^{5}$ correspond to ECM trained with a batch size of $512$ for $200$k, $400$k, and $800$k iterations, respectively, with the smoothing factor $c=0.03$ enabled in the adaptive weighting $w(\Delta)$. We use $\bar{w}(t)=\nicefrac{{1}}{{t}}$ as the timestep weighting function to train all these models as this $\bar{w}(t)$ achieves good performance on $\text{FD}_{\text{DINOv2}}$.
Scaling of Model Size and Model FLOPs.
We include details of model capacity as well as FLOPs in Tab. 8 to replicate this plot on ImageNet 64$\times$64.
On ImageNet 64$\times$64, we scale up the training budgets of ECM-S and ECM-XL from $12.8\mathrm{M}$ (batch size of $128$ and $100$k iterations) to $102.4\mathrm{M}$ (batch size of $1024$ and $100$k iterations). We empirically find that scaling the base learning rate by $\sqrt{n}$ works well when scaling the batch size by a factor of $n$ when using Adam (Kingma and Ba, 2014) optimizer.